OpenStudy (anonymous):

find (fog)(2) and (f+g)(2) when f(x)=1/x and g(x)=4x+9

OpenStudy (anonymous):

fog(2) = f(g(2)) f(x) = 1/x g(x) = 4x + 9 So f(g(2)) = 1/4x2 + 9 1/17 Am I right?

OpenStudy (anonymous):

That's what I got for fog yes but what's (f+g)(2) then

OpenStudy (anonymous):

@kennybm

OpenStudy (anonymous):

eeermmm wait wait I'm not really sure but lets try f(2) and g(2) for it?

OpenStudy (anonymous):

so its 17.5?

OpenStudy (anonymous):

waiiit @ganeshie8 how do we do (f+g)(2)

ganeshie8 (ganeshie8):

(f+g)(2) = f(2) + g(2)

ganeshie8 (ganeshie8):

(f+g)(x) = f(x) + g(x) we just add both the functions

OpenStudy (anonymous):

so then it is 17.5 @ganeshie8 ?

OpenStudy (anonymous):

1/2 + 17

OpenStudy (anonymous):

yeaaah I guess that's right.